what is the difference between
3(x –1) + 4 – 2x
3(x-1) +4 = 2x

Answer by praseenakos@yahoo.com(507)

QUESTION:
what is the difference between
3(x –1) + 4 – 2x
3(x-1) +4 = 2x


ANSWER:

3(x –1) + 4 – 2x is an expression

and 3(x-1) +4 = 2x is an equation.


An expression is a combination of numbers, operators, grouping symbols (such as brackets and parentheses) and/or free variables and bound variables arranged in a meaningful way which can be evaluated.




An expression which includes an equal sign is called an equation.


Otherwise we can define,


An equation is a mathematical statement, in symbols, that two things are the

same (or equivalent). Equations are written with an equal sign, as in
2 + 3 = 5.






Hope you understood.


Regards.



Praseena.